QUESTION 3 Using the data below, determine the bending factor of safety for the pinion. Gear type: Spur Gear with Full depth teeth . Pressure angle (degrees) 20 . Np (teeth) 20 * n (rpms) 1200 ·H(hp) = 60 . Ng (teeth) 36 Pd (teeth/in) 4 face width (in) 3 . Quality factor (Q ) = 8 e gear mounting straddled between bearings Pinion (HB) 300 (material: steel) Gear (HB) = 240 (material: steel) . Are the teeth crowned? No Type of gear box= Precision and enclosed . Loading (source, output)-Uniform-Moderate shock . Cycles of the pinion (cycles) 1047 . Reliability (R) 0.99 . Assume that the gear was solid and no rim was made . Pinion and gear material grade 2 through hardened steel . Assume gears were not altered at assembly (Ce 1)
Answer #1

No Given ressure angle= p 20 -Te-36 20 4-4 test in) Pacero ic h=6-3 n: 76.2 mm module= Ypd= 1/4 = 0,25 fn/ teeth- 6-35 mm HP-2 2 No b TL.m feニ47786.56 N rO1 Cu Dia.of pinion = pf= m.Tpこ6.35χ20= !27 mm Nele city of pinion : Th DN 12구지2cD 60000 60000 =